# Higher Order Dualities between Prime Ideals

### v1

@@ 4 A New Formula for the Chebotarev Density in arbitrary finite Galois extensions
$\sum_{2\leq N(I)\leq X}Q^{2}(I)=$ $\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\int_{Y}^{N(\mathfrak{q})}\Psi^{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi^{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}$
$+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\right)+O\left(\frac{x}{e^{c_{3}\sqrt{\log X\log\log X}}}\right)$ (4.19)
@@
$\sum_{2\leq N(I)\leq X}Q^{2}(I)=c_{K}\cdot X+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\right)$ (4.21)
@@
$\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\int_{Y}^{N(\mathfrak{q})}\Psi^{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi^{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}$
$=c_{K}\cdot X+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log X\log\log X}{e^{c_{4}\sqrt{\log Y}}}\right)$ (4.22)
Interestingly, the LHS of (4.22) is exactly the expression in brackets in the RHS of (4.12) that we were required to estimate. Therefore, now combining (4.1), (4.12) and (4.22), we finally get that
$\sum_{2\leq N(I)\leq X}Q_{C}^{2}(I)=c_{K}\cdot\frac{|C|}{|G|}\cdot X+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log X\log\log X}{e^{c_{4}\sqrt{\log Y}}}\right)$ (4.23)
@@
Recalling Theorem 1.1 from [SW18], we see that it is also an asymptotic relation, and also, surprisingly has the same right hand side. Precisely,
$\sum_{2\leq N(I)\leq X}Q_{C}(I)\sim c_{K}\cdot\frac{|C|}{|G|}\cdot X$ (4.25)
It is indeed intriguing to see that irrespective of the $k^{th}$ largest norm that we choose, the sum on the left has the same asymptote due to the use of the Chebotarev Density Theorem. One can also prove the same asymptotic result for other $k^{th}$ largest norms, as done in the classical case for any arbitrary $k^{th}$ largest prime factors (see [AS(ip)]). We now prove our main theorem that provides us the formula for the Chebotarev Density using the first and second order duality.
@@
$\lim_{X\to\infty}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\frac{\mu_{K}(I)(\omega_{K}(I)-1)}{N(I)}=\lim_{X\to\infty}\sum_{2\leq N(I)\leq X}\frac{\mu_{K}(I)(\omega_{K}(I)-1)f(I)}{N(I)}=\frac{|C|}{|G|}$
@@
$=\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\left[c_{k}\cdot\frac{X}{N(I)}\right]$ (4.26)
@@
$\#\{J\subset\mathcal{O}_{K}:J\subset I;\;N(J)\leq X\}=\sum_{\begin{subarray}{c}J\subset\mathcal{O}_{K}\\ N(J)\leq X\\ I\supset J\end{subarray}}1=c_{K}\cdot\frac{X}{N(I)}+O\left(\left(\frac{X}{N(I)}\right)^{1-\frac{1}{d}}\right)$ (4.27)
@@
$=\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\sum_{\begin{subarray}{c}J\subset\mathcal{O}_{K}\\ N(J)\leq X\\ I\supset J\end{subarray}}1+O\left(\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\left(\frac{X}{N(I)}\right)^{1-\frac{1}{d}}\right)$ (4.28)
Let us first estimate the error term on the right of (4.28). To do so, we recall the following bound of the sum of the generalized Möbius function (see Lemma 2.2 (ii) in [SW18]) that follow analogously from their classical counterparts:
$\sum_{N(I)\leq X}\frac{\mu(I)}{N(I)}$ $=O\left(e^{-c_{6}(\log X)^{\frac{1}{12}}}\right)$ (4.29)
@@
$\ll\frac{X\log\log X}{e^{c_{6}(\log X)^{\frac{1}{12}}}}$ (4.31)
Now, for the main term in the RHS of (4.28), we observe with a interchange in summation that
$\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\sum_{\begin{subarray}{c}J\subset\mathcal{O}_{K}\\ N(J)\leq X\\ I\supset J\end{subarray}}1$ $=\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{I\supset J}\mu_{K}(I)\omega_{K}(I)f(I)$ (4.32)
@@
$\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)\omega_{K}(I)f(I)=\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)(\omega_{K}(I)-1)f(I)+\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)f(I)$ (4.33)
@@
$\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)\omega_{K}(I)f(I)=\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)(\omega_{K}(I)-1)f(I)+\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)f(I)$ (4.34)
@@
$=\frac{|C|}{|G|}\cdot X+O\left(\frac{X(\log\log X)^{2}}{\log X}\right)$ (4.35)
Further, applying Lemma 2.2 and then Theorem 1.1 of [SW18] in the second in the RHS of (4.35), we ave
$\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)f(I)$ $=-\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}Q_{C}^{1}(I)$
$=-\frac{|C|}{|G|}\cdot X+O\left(\frac{X}{e^{c_{5}(\log X)^{\frac{1}{3}}}}\right)$ (4.36)
Therefore, combining (4.32), (4.34), (4.35) and (4.36), we get
$\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\sum_{\begin{subarray}{c}J\subset\mathcal{O}_{K}\\ N(J)\leq X\\ I\supset J\end{subarray}}1\ll\frac{X(\log\log X)^{2}}{\log X}$ (4.37)
@@
$X\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}\ll\frac{X(\log\log X)^{2}}{(\log X)}$ (4.38)
@@
$\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}\ll\frac{(\log\log X)^{2}}{(\log X)}$ (4.39)
@@
$=\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\frac{\mu_{K}(I)(\omega_{K}(I)-1)}{N(I)}-\frac{|C|}{|G|}+O\left(e^{-c_{6}(\log X)^{\frac{1}{12}}}\right)$ (4.40)
@@
$\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\frac{\mu_{K}(I)(\omega_{K}(I)-1)}{N(I)}=\frac{|C|}{|G|}+O\left(\frac{(\log\log X)^{2}}{\log X}\right)$ (4.41)
@@
It is important to note here that this method to obtain the formula is significantly different from any that has been used earlier. The method used in [SW18] uses the sums of $\mu_{K}(I)\log N(I)$, which faces difficulties in our case as the error estimates grow large due to the largeness of $\Psi_{K,2}(X,Y)$. Difficulties were also faced in following a method similar to the one used in [Se25] due to absence of a one-one correspondence between the multiples of an ideal $I\subset\mathcal{O}_{K}$ and multiples of its norm $N(I)$.
@@ 5 Two Estimates of Sums involving $\mu_{K}(I)$ and $\omega_{K}(I)$
In [Al77], [AJ24] and [Se25], we observe that the authors have used the estimates of sums involving $\mu(n)$ and $\omega(n)$ which then are required to prove the respective final results. As can be observed in §4, such estimates were not required to obtain our main result. Yet, we still are capable to find estimates of such sums with the machinery that we have developed. In this section, therefore, we will prove two theorem that give us quantitative estimates of such sums, which although are not necessary to get the formula for the Chebotarev Density, can be viewed as an application of our duality identities.
@@
$\sum_{\begin{subarray}{c}N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\ll\frac{X\log X}{e^{c_{10}(\log X)^{\frac{1}{12}}}}$
@@
$\sum_{\begin{subarray}{c}N(I)\leq X\\ Ic\in S(L/K;C)\end{subarray}}\mu_{K}(I)=\sum_{N(I)\leq X}\mu_{K}(I)f(I)$ (5.1)
@@
$\sum_{N(I)\leq X}\mu_{K}(I)f(I)=-\sum_{N(I)\leq X}\sum_{JJ^{\prime}=I}\mu_{K}(J^{\prime})Q_{C}^{1}(J)$ (5.2)
Note, the second sum runs over all ideals $J\supset I$ and we write $Q_{C}(J)=Q_{C}^{1}(J)$. Truncating the double sum at $\sqrt{X}$ and using the hyperbola method, we deduce
$\sum_{N(I)\leq X}\mu_{K}(I)f(I)=-\sum_{N(J^{\prime})\leq\sqrt{X}}\mu_{K}(J^{\prime})\sum_{N(J)\leq\frac{X}{N(J^{\prime})}}Q_{C}^{1}(J)-\sum_{N(J)\leq\sqrt{X}}Q_{C}^{1}(J)\sum_{\sqrt{X}<N(J^{\prime})\leq\frac{X}{N(J)}}\mu_{K}(J^{\prime})$ (5.3)
@@
$\sum_{N(J^{\prime})\leq\sqrt{X}}\mu_{K}(J^{\prime})\sum_{N(J)\leq\frac{X}{N(J^{\prime})}}Q_{C}^{1}(J)=\sum_{N(J^{\prime})\leq\sqrt{X}}\mu_{K}(J^{\prime})\left[c_{K}\cdot\frac{|C|}{|G|}\cdot\frac{X}{N(J^{\prime})}+O\left(\frac{X}{N(J^{\prime})e^{c_{5}\left(\log\frac{X}{N(J^{\prime})}\right)^{\frac{1}{3}}}}\right)\right]$ (5.4)
@@
$\sum_{N(I)\leq X}\mu_{K}(I)$ $=O\left(Xe^{-c_{6}(\log X)^{\frac{1}{12}}}\right)$ (5.6)
@@
$\sum_{N(J)\leq\sqrt{X}}Q_{C}^{1}(J)\sum_{\sqrt{X}<N(J^{\prime})\leq\frac{X}{N(J)}}\mu_{K}(J^{\prime})\ll\sum_{N(J)\leq\sqrt{X}}\frac{X}{N(J)e^{c_{6}\left(\log\frac{X}{N(J)}\right)^{\frac{1}{12}}}}\ll\frac{X\log X}{e^{c_{9}(\log X)^{\frac{1}{12}}}}$ (5.7)
Here, we use the fact that $Q_{C}^{1}(I)$ is bounded by the degree of extension $K/\mathbb{Q}$. Therefore, combining (5.3), (5.5) and (5.7), and choosing $c_{10}=\min\{c_{8},c_{9}\}$, we have
$\sum_{N(I)\leq X}\mu_{K}(I)f(I)\ll\frac{X\log X}{e^{c_{10}(\log X)^{\frac{1}{12}}}}$ (5.8)
@@
We observe that, here we used the first order duality lemma. The proofs in [SW18] did not require such an estimate and therefore, Theorem 5.1 was not proved then.
@@
$\sum_{\begin{subarray}{c}N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\ll\frac{X(\log\log X)^{4}}{\log X}$
@@
$\sum_{N(I)\leq X}\mu_{K}(I)(\omega_{K}(I)-1)f(I)=\sum_{N(I)\leq X}\sum_{\begin{subarray}{c}JJ^{\prime}=I\\ Q^{1}(J)=1\end{subarray}}\mu_{K}(J^{\prime})Q_{C}^{2}(J)$ (5.10)
@@
$\sum_{N(I)\leq X}\mu_{K}(I)(\omega_{K}(I)-1)f(I)=\sum_{N(J^{\prime})\leq T}\mu_{K}(J^{\prime})\sum_{\begin{subarray}{c}N(J)\leq\frac{X}{N(J^{\prime})}\\ Q^{1}(J)=1\end{subarray}}Q_{C}^{2}(J)+\sum_{\begin{subarray}{c}N(J)\leq\frac{X}{T}\\ Q^{1}(J)=1\end{subarray}}Q_{C}^{2}(J)\sum_{T<N(J^{\prime})\leq\frac{X}{N(J)}}\mu_{K}(J^{\prime})$ (5.11)
@@
$\ll\frac{X}{e^{c_{6}(\log T)^{\frac{1}{12}}}}+\frac{X\log T(\log\log X)^{2}}{\log\left(\frac{X}{T}\right)}$ (5.12)

### v2

@@ 4 A New Formula for the Chebotarev Density in arbitrary finite Galois extensions
$\sum_{2\leq N(I)\leq X}Q^{2}(I)=$ $\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}$
$+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\right)+O\left(\frac{x}{e^{c_{3}\sqrt{\log X\log\log X}}}\right).$ (4.19)
@@
$\sum_{2\leq N(I)\leq X}Q^{2}(I)=c_{K}\cdot X+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\right).$ (4.21)
@@
$\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}$
$=c_{K}\cdot X+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log X\log\log X}{e^{c_{4}\sqrt{\log Y}}}\right).$ (4.22)
Interestingly, the LHS of (4.22) is exactly the expression in brackets in the RHS of (4.12) that we were required to estimate. Therefore, now combining (4.1), (4.12), and (4.22), we finally get that
$\sum_{2\leq N(I)\leq X}Q_{C}^{2}(I)=c_{K}\cdot\frac{|C|}{|G|}\cdot X+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log X\log\log X}{e^{c_{4}\sqrt{\log Y}}}\right).$ (4.23)
@@
Recalling Theorem 1.1 from [SW18], we see that it is also an asymptotic relation, and also, surprisingly, has the same right hand side. Precisely,
$\sum_{2\leq N(I)\leq X}Q_{C}(I)\sim c_{K}\cdot\frac{|C|}{|G|}\cdot X.$ (4.25)
It is indeed intriguing to see that, irrespective of the $k^{th}$ largest norm that we choose, the sum on the left has the same asymptote due to the use of the Chebotarev Density Theorem. One can also prove the same asymptotic result for other $k^{th}$ largest norms, as done in the classical case for any arbitrary $k^{th}$ largest prime factors (see [AS(ip)]). We now prove the main theorem that provides us with the formula for the Chebotarev Density using the first and second order duality.
@@
$\lim_{X\to\infty}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\frac{\mu_{K}(I)(\omega_{K}(I)-1)}{N(I)}=\lim_{X\to\infty}\sum_{2\leq N(I)\leq X}\frac{\mu_{K}(I)(\omega_{K}(I)-1)f(I)}{N(I)}=\frac{|C|}{|G|}.$
@@
$=\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\left[c_{K}\cdot\frac{X}{N(I)}\right].$ (4.26)
@@
$\#\{J\subset\mathcal{O}_{K}:J\subset I;\;N(J)\leq X\}=\sum_{\begin{subarray}{c}J\subset\mathcal{O}_{K}\\ N(J)\leq X\\ I\supset J\end{subarray}}1=c_{K}\cdot\frac{X}{N(I)}+O\left(\left(\frac{X}{N(I)}\right)^{1-\frac{1}{d}}\right).$ (4.27)
@@
$=\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\sum_{\begin{subarray}{c}J\subset\mathcal{O}_{K}\\ N(J)\leq X\\ I\supset J\end{subarray}}1+O\left(\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\left(\frac{X}{N(I)}\right)^{1-\frac{1}{d}}\right).$ (4.28)
Let us first estimate the error term on the right of (4.28). To do so, we recall the following bound of the sum of the generalized Möbius function (see Lemma 2.2 (ii) in [SW18]) that follows analogously from their classical counterparts:
$\sum_{N(I)\leq X}\frac{\mu_{K}(I)}{N(I)}$ $=O\left(e^{-c_{6}(\log X)^{\frac{1}{12}}}\right).$ (4.29)
@@
$\ll\frac{X\log\log X}{e^{c_{6}(\log X)^{\frac{1}{12}}}}.$ (4.31)
Now, for the main term in the RHS of (4.28), we observe with an interchange in summation that
$\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\sum_{\begin{subarray}{c}J\subset\mathcal{O}_{K}\\ N(J)\leq X\\ I\supset J\end{subarray}}1$ $=\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{I\supset J}\mu_{K}(I)\omega_{K}(I)f(I).$ (4.32)
@@
$\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)\omega_{K}(I)f(I)=\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)(\omega_{K}(I)-1)f(I)+\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)f(I).$ (4.33)
@@
$\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)\omega_{K}(I)f(I)=\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)(\omega_{K}(I)-1)f(I)+\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)f(I).$ (4.34)
@@
$=\frac{|C|}{|G|}\cdot X+O\left(\frac{X(\log\log X)^{2}}{\log X}\right).$ (4.35)
Further, applying Lemma 2.2 and then Theorem 1.1 of [SW18] in the second sum in the RHS of (4.34), we have
$\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)f(I)$ $=-\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}Q_{C}^{1}(J)$
$=-\frac{|C|}{|G|}\cdot X+O\left(\frac{X}{e^{c_{5}(\log X)^{\frac{1}{3}}}}\right).$ (4.36)
Therefore, combining (4.32), (4.34), (4.35), and (4.36), we get
$\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\sum_{\begin{subarray}{c}J\subset\mathcal{O}_{K}\\ N(J)\leq X\\ I\supset J\end{subarray}}1\ll\frac{X(\log\log X)^{2}}{\log X}.$ (4.37)
@@
$X\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}\ll\frac{X(\log\log X)^{2}}{(\log X)}.$ (4.38)
@@
$\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}\ll\frac{(\log\log X)^{2}}{(\log X)}.$ (4.39)
@@
$=\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\frac{\mu_{K}(I)(\omega_{K}(I)-1)}{N(I)}-\frac{|C|}{|G|}+O\left(e^{-c_{6}(\log X)^{\frac{1}{12}}}\right).$ (4.40)
@@
$\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\frac{\mu_{K}(I)(\omega_{K}(I)-1)}{N(I)}=\frac{|C|}{|G|}+O\left(\frac{(\log\log X)^{2}}{\log X}\right).$ (4.41)
@@
It is important to note here that this method to obtain the formula is significantly different from any that has been used earlier. The method used in [SW18] uses the sums of $\mu_{K}(I)\log N(I)$, which faces difficulties in our case as the error estimates grow large due to the largeness of $\Psi_{K,2}(X,Y)$. Difficulties were also faced in following a method similar to the one used in [Se25] due to the absence of a one-to-one correspondence between the multiples of an ideal $I\subset\mathcal{O}_{K}$ and multiples of its norm $N(I)$.
@@ 5 Two Estimates of Sums involving $\mu_{K}(I)$ and $\omega_{K}(I)$
In [Al77], [AJ24] and [Se25], we observe that the authors have used the estimates of sums involving $\mu(n)$ and $\omega(n)$, which are required to prove the respective final results. As can be observed in §4, such estimates were not required to obtain our main result. Yet, we are capable of finding estimates of such sums with the machinery that we have developed. In this section, therefore, we will prove two theorems that give us quantitative estimates of such sums, which, although they are not necessary to get the formula for the Chebotarev Density, can be viewed as an application of our duality identities.
@@
$\sum_{\begin{subarray}{c}N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\ll\frac{X\log X}{e^{c_{10}(\log X)^{\frac{1}{12}}}}.$
@@
$\sum_{\begin{subarray}{c}N(I)\leq X\\ Ic\in S(L/K;C)\end{subarray}}\mu_{K}(I)=\sum_{N(I)\leq X}\mu_{K}(I)f(I).$ (5.1)
@@
$\sum_{N(I)\leq X}\mu_{K}(I)f(I)=-\sum_{N(I)\leq X}\sum_{JJ^{\prime}=I}\mu_{K}(J^{\prime})Q_{C}^{1}(J).$ (5.2)
Note, the second sum runs over all ideals $J\supset I$, and we write $Q_{C}(J)=Q_{C}^{1}(J)$. Truncating the double sum at $\sqrt{X}$ and using the hyperbola method, we deduce
$\sum_{N(I)\leq X}\mu_{K}(I)f(I)=-\sum_{N(J^{\prime})\leq\sqrt{X}}\mu_{K}(J^{\prime})\sum_{N(J)\leq\frac{X}{N(J^{\prime})}}Q_{C}^{1}(J)-\sum_{N(J)\leq\sqrt{X}}Q_{C}^{1}(J)\sum_{\sqrt{X}<N(J^{\prime})\leq\frac{X}{N(J)}}\mu_{K}(J^{\prime}).$ (5.3)
@@
$\sum_{N(J^{\prime})\leq\sqrt{X}}\mu_{K}(J^{\prime})\sum_{N(J)\leq\frac{X}{N(J^{\prime})}}Q_{C}^{1}(J)=\sum_{N(J^{\prime})\leq\sqrt{X}}\mu_{K}(J^{\prime})\left[c_{K}\cdot\frac{|C|}{|G|}\cdot\frac{X}{N(J^{\prime})}+O\left(\frac{X}{N(J^{\prime})e^{c_{5}\left(\log\frac{X}{N(J^{\prime})}\right)^{\frac{1}{3}}}}\right)\right].$ (5.4)
@@
$\sum_{N(I)\leq X}\mu_{K}(I)$ $=O\left(Xe^{-c_{6}(\log X)^{\frac{1}{12}}}\right).$ (5.6)
@@
$\sum_{N(J)\leq\sqrt{X}}Q_{C}^{1}(J)\sum_{\sqrt{X}<N(J^{\prime})\leq\frac{X}{N(J)}}\mu_{K}(J^{\prime})\ll\sum_{N(J)\leq\sqrt{X}}\frac{X}{N(J)e^{c_{6}\left(\log\frac{X}{N(J)}\right)^{\frac{1}{12}}}}\ll\frac{X\log X}{e^{c_{9}(\log X)^{\frac{1}{12}}}}.$ (5.7)
Here, we use the fact that $Q_{C}^{1}(I)$ is bounded by the degree of extension $K/\mathbb{Q}$. Therefore, combining (5.3), (5.5), and (5.7), and choosing $c_{10}=\min\{c_{8},c_{9}\}$, we have
$\sum_{N(I)\leq X}\mu_{K}(I)f(I)\ll\frac{X\log X}{e^{c_{10}(\log X)^{\frac{1}{12}}}}.$ (5.8)
@@
We observe that here we used the first order duality lemma. The proofs in [SW18] did not require such an estimate and therefore, Theorem 5.1 was not proved then.
@@
$\sum_{\begin{subarray}{c}N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\ll\frac{X(\log\log X)^{4}}{\log X}.$
@@
$\sum_{N(I)\leq X}\mu_{K}(I)(\omega_{K}(I)-1)f(I)=\sum_{N(I)\leq X}\sum_{\begin{subarray}{c}JJ^{\prime}=I\\ Q^{1}(J)=1\end{subarray}}\mu_{K}(J^{\prime})Q_{C}^{2}(J).$ (5.10)
@@
$\sum_{N(I)\leq X}\mu_{K}(I)(\omega_{K}(I)-1)f(I)=\sum_{N(J^{\prime})\leq T}\mu_{K}(J^{\prime})\sum_{\begin{subarray}{c}N(J)\leq\frac{X}{N(J^{\prime})}\\ Q^{1}(J)=1\end{subarray}}Q_{C}^{2}(J)+\sum_{\begin{subarray}{c}N(J)\leq\frac{X}{T}\\ Q^{1}(J)=1\end{subarray}}Q_{C}^{2}(J)\sum_{T<N(J^{\prime})\leq\frac{X}{N(J)}}\mu_{K}(J^{\prime}).$ (5.11)
@@
$\ll\frac{X}{e^{c_{6}(\log T)^{\frac{1}{12}}}}+\frac{X\log T(\log\log X)^{2}}{\log\left(\frac{X}{T}\right)}.$ (5.12)